OpenStudy (sphott51):

What is the approximate solution of 3^x-1 = 8.2?

OpenStudy (flcon04):

Any choices?

OpenStudy (anonymous):

is it \[\huge 3^{x-1}=8.2\]?

OpenStudy (flcon04):

x=3.06

OpenStudy (anonymous):

unlikely

OpenStudy (sphott51):

Yes @satellite73

OpenStudy (tkhunny):

\(3^{1} = 3\) \(3^{2} = 9\) The question is no good. "The approximation" doesn't exist. "AN" approximation can be generated. Linear Interpolation will produce AN approximation: \(1 + \dfrac{8.2-3}{9-3} = 1.8667\), Thus "AN" approximation for x is 2.8667. There are infinitely many approximations. One might also observe, \(3^{x-1} = \dfrac{3^{x}}{3} = 8.2 \implies 3^{x} = 24.6\) \(3^{3} = 27\) And we have another Linear Interpolation will produce AN approximation: \(2 + \dfrac{24.6-9}{27-9} = 2.8667\), Thus "AN" approximation for x is 2.8667. It should be no surprise that these produce the same result. Any other ideas? Calculus? Iteration of some sort?

OpenStudy (anonymous):

Refer to the attachment.